# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ V-A2 DL-based WF Attack models
Singh et al. [141] presented AutoKeras, a neural architecture search (NAS) tool, to explore whether automated DNN architecture design can improve WF attacks compared to manually crafted models. While NAS-generated models require larger exploration budgets and longer inference times, they highlight the potential for automated deep-learning-driven WF attacks. Zhuo et al. [70] introduced WF-Transformer, the first WF attack model based on Transformer networks, which is designed to capture the long-term dependencies in traffic traces, a feature that traditional CNN-based models fail to extract. The model uses packet direction as input and employs a sophisticated design of the Transformer network, including relative position embedding and multi-head attention mechanisms, to improve feature extraction. Jansen et al. [154] investigated the real-world feasibility of WF attacks on Tor by addressing a critical gap in previous research: the mismatch between training data collected from Tor exit relays and real-world WF attacks conducted from entry relays. To bridge this gap, the authors introduce Retracer, which transforms exit-side traces into entry-side traces, improving classifier training and yielding more accurate real-world WF threat estimates.
@@
To address the limitation associated with WF attacks and defense strategies in the presence of noise, Huang et al. [155] proposed a filter-assisted attack (FAA) to counter random packet defenses (RPD) and introduced a list-assisted defense (LAD) to mitigate FAA, balancing defense efficacy and overhead. The next year, Qian et al. [123] investigated the impact of noise on WF attacks and introduced two attack methods: filter-assisted attack and augmentation-assisted attack. The first attack method leverages the packet size distribution to effectively filter out noise, while the second one trains a classification model by incorporating artificial noise. Experimental results show the efficacy of defending against different attacks and noises. Bahramali et al. [71] identified that WF classifiers often fail when deployed in unobserved network conditions due to variations in bandwidth, Tor circuits, and traffic patterns. To overcome this limitation, they proposed a network trace augmentation technique, NetAugment, to enhance the robustness of WF models. The NetAugment was instantiated through semi-supervised and self-supervised learning techniques.
@@
Through analyzing the present WF attacks, Mathews et al. [30] found that a multi-channel input format that provides richer contextual information and enables the model to learn robust representations even in the presence of heavy traffic obfuscation. Then, the authors introduced Laserbeak, a state-of-the-art WF attack designed to break Tor defenses by leveraging multi-channel feature representation and transformer-based deep learning models. LASERBEAK demonstrates absolute performance improvements of up to 36.2% compared with prior attacks against defended traffic.
@@
The deep learning based WF attacks often ignore low-level timing information and use a simple direction-based representation method that only retains the spatial information of website traffic packets. However, the timing information plays a significant role in the WF attack under the research of Hong et al. [38]. They presented a WF attack technique called Time-Sampling, which leverages both spatial and temporal timing information from website traffic to improve the accuracy of WF attacks. The Time-Sampling model uses a time-based sampling technique that fills packet gaps with zeros, preserving both timing and spatial information in the traffic. Through evaluation, the accuracy is improved by 20% at most, higher than direction-based attacks.
@@
Traditional WF techniques either rely on single-flow features, which are sensitive to network changes, or multi-flow aggregation methods that fail to capture temporal and spatial relationships between network flows. To solve this issue, Tan et al. [75] presented STC-WF, a WF attack that leverages spatio-temporal correlations among network flows to improve classification accuracy, particularly in encrypted traffic scenarios. The STC-WF uses a Graph Neural Network (GNN)-based approach that models these correlations as a Spatio-Temporal Correlation Graph (STCG). To overcome the limitation of traditional WF techniques, which struggle with content updates and labeling overhead, limiting their effectiveness in real-world scenarios. Gao et al. [31] introduced MRCGCN, a self-supervised learning-based WF attack that leverages multi-level resource-coherent GNNs to improve the accuracy and robustness of WF attacks. MRCGCN overcomes these challenges by representing website traffic as structured graphs at resource, flow, and host levels and applying graph learning techniques for classification.
@@
Concept drift in website fingerprinting can occur when websites modify their structure or content. Even small changes, such as adding images, updating ads, or reorganizing page elements, alter the resulting traffic patterns. These shifts cause the distribution of features like packet sizes, timing, and traffic length to deviate from the data on which models were trained. To address the problem of concept drift, Ding et al. [44] presented LRCT, a DL-based WF attack model designed to improve classification accuracy and efficiency while addressing concept drift—a phenomenon where model performance degrades over time due to network changes. The LRCT network effectively leverages the temporal learning advantages of Local RNN and the spatial learning strengths of CNN by designing the local feature extraction block (denoted as LRC Block). The network then uses a Transformer Encoder to capture more robust global features from the low-dimensional data. Experimental results comparing LRCT with state-of-the-art WF methods demonstrate that the proposed LRCT achieves 99.34% accuracy in a closed-world scenario, outperforming other benchmark models.
@@
Ginige et al. [156] proposed TrafficGPT, the first solution that leverages GPT-2 for feature extraction in encrypted traffic classification. GPT-2’s ability to understand sequential data is utilized to extract comprehensive features, improving the performance of open-set classification. This paper focuses on the open-set scenario, where the model must distinguish between known and unknown classes. This is particularly important in real-world applications, such as detecting harmful content while allowing benign traffic. Yuan et al. [147] introduced Class Incremental Fingerprinting (CIF), a WF attack designed to continuously learn and expand its classification capabilities in real-world dynamic environments. Traditional DL-based WF attacks achieve high accuracy in static experimental settings but struggle in open-world scenarios, where new websites constantly emerge. CIF addresses this limitation by incorporating Class Incremental Learning (CIL) techniques, allowing the attack model to adapt to new monitored websites without forgetting previously learned knowledge. Yang et al. [146] introduced DTPN (Diffusion-based Traffic Purification Network), which counteract adversarial perturbations in Tor WF defenses by leveraging diffusion models. Unlike traditional adversarial training, which struggles with generalization and frequent updates, DTPN purifies protected network traffic by removing adversarial noise, enabling existing WF classifiers to maintain high accuracy without retraining. The proposed UNetFormer architecture, combining TransformerBlocks and one-dimensional convolutions, estimates noise effectively, while a custom loss function enhances model generalization. Mei et al. [145] presented HPETC, a High Precision and Efficient Traffic Classification system designed to classify Tor traffic in real-world, large-scale network environments. Unlike previous Traffic Classification (TC) and WF techniques, HPETC effectively mitigates the base rate fallacy, a common issue where false positives overwhelm the classification process due to the extremely low proportion of anonymous traffic compared to normal network traffic.
@@ V-A3 Multi-tab WF attacks
BAPM [117],
Walkie-Talkie [45]
@@
ARES [109] 91.30 91.30 90.20 81.80 75.20 58.70 52.80 43.80
FingerMamba [126]
@@
MTWFAs [13] FCN
DS-19 [57],
ARES [109]
@@
FMWF [132] Few-shotFine-tuning
ARES [109]
@@
ARES [109],
Deng et al. [32],
@@
ARES [109] and
TMWF [43]
@@
Traditional WF attacks are performed on a well-constructed dataset, in which each packet trace is generated by visiting only one page tab at a time. Thus, attackers only need to analyze traces entirely belonging to a specific website without any noise data, which is known as the single-tab assumption. This assumption makes WF attacks less difficult, but it is unrealistic: users may engage in other activities, such as instant messaging alongside web browsing, and similar behavior is entirely random  [37]. In addition, even if users only browse the web, they may open a new page tab before the last one finishes loading, a behavior known as multi-tab browsing [80, 157, 6], such as jumping to another page through a hyperlink. A comparison of representative multi-tab WF attacks with 10% overlapping proportions as shown in TABLE IX.
@@
After Juarez et al. [6] pointed out that multi-tab browsing behavior significantly degrades the performance of WF models under the single-tab assumption, Wang and Goldberg [37] attempted to transform multi-tab traces into single-tab traces by splitting packet sequences and processing them using existing single-tab WF classifiers. In  [37], multi-tab traffic traces are categorized into three fundamental types based on their time interval relationships: Positive-time, Zero-time, and Negative-time. Figure 6 illustrates these scenarios for 2-tab traffic traces: (1) Positive-time separated traces refer to instances where the user lingers on a web page for a period before navigating to the next one, resulting in a brief lull in network activity and a noticeable gap in traffic. This time gap is significant enough to be detected as a pause but still shorter than the threshold used in time-based traffic splitting. (2) Zero-time separated traces refer to scenarios where a user clicks a link on a web page that is still loading, interrupting the current page and immediately initiating requests for the next one. This results in a clear transition between two web pages without any noticeable time gap between them. (3) Negative-time separated traces refer to situations where the user is loading two web pages simultaneously in multiple tabs. In such cases, the appropriate split point is considered to be the moment the second page begins loading. This is the most challenging class to separate, as there is neither a noticeable time gap nor a distinct cell pattern signaling the transition. Nonetheless, machine learning techniques can still be used to identify the split by extracting informative features from the cell sequences.
@@
Wang and Goldberg [37] focused on identifying split points within multi-tab traffic to divide it into multiple different single-tab segments for classification. However, this segmentation process results in information loss in the overlapping portions of the traffic. Although the algorithm improves the accuracy of split-point detection, it can only predict the first webpage visited in a multi-tab browsing session. Moreover, as the number of tabs increases, accurately locating the split points becomes increasingly difficult [119]. Xu et al. [62] introduced the BalanceCascade-XGBoost algorithm to identify the split point between two tabs and classify traffic according to the initial chunk of data. Later, Yin et al. [65] improved the BalanceCascade-XGBoost algorithm by applying Split Point Identification and Chunk-Based Classification mechanism. Cui et al. [80] utilized the Hidden Markov Model to segment continuous browsing traffic and conducted webpage recognition on each segment through K-NN, achieving an impressive 80% accuracy in identifying the split points. To determine the final category of the webpage, a majority vote was employed. However, it treats overlapping areas equally with other areas and fails to consider the relevance and importance of the overlapping area in determining the webpage category [143].
@@
In 2020, Cui et al. [120] utilized a CNN model to differentiate between single-tab and multi-tab traces, and extracted a fixed number of packets from the beginning and end of 2-tab traces. They then used CNN, Long Short-Term Memory (LSTM), and Stacked Denoising Auto-Encoding (SDAE) models to classify the extracted pure trace segments, further improving the accuracy of WF. However, their approach is still constrained by a limited number of page configurations and relies on multi-tab traces containing sufficiently long and non-overlapping pure segments. Then, Guan et al. [117] proposed a block attention profiling model (BAPM), which focuses on overlapping webpage traffic. However, with the number of webpages increasing, BAPM faces the challenge of dynamically adapting the number of attention heads, which introduces instability to the network structure and leads to more parameters. In addition, BAPM only considers a single scenario of overlapping visits and cannot adapt to more complex browsing scenarios. Previous CNN-based WF attacks rely on a carefully curated dataset where each packet trace is created by loading only one webpage at a time. Under this single-tab assumption, attackers analyze clean, noise-free traces, each belonging exclusively to a single website. However, the assumptions of previous work are obviously unreal. Users may access multiple webpages in multiple tabs at the same time.
@@
Guan et al. [117] found that the existing multi-tab attacks fail to exploit the overlapping regions, which contain mixed data from adjacent page tabs. To address this limitation, they leveraged the CNN to generate a tab-aware representation on the whole packet trace, including overlapping areas, and split the representation into blocks to reduce the influence of mixed data under the traffic direction features. Using multi-head attention, the model groups blocks belonging to the same website, allowing it to identify multiple websites simultaneously from a global perspective. This approach eliminates the need for explicit feature engineering and adapts naturally to different degrees of overlap.
@@
Later, Jin et al. [43] introduced Transformer-based Multi-tab Website Fingerprinting (TMWF), an end-to-end deep learning model that leverages a Transformer architecture to improve multi-tab WF attacks. Unlike previous models, TMWF does not require prior knowledge of the number of tabs in a browsing session. It treats multi-tab WF recognition as a set prediction problem, similar to object detection models. By introducing queries for tabs to extract WF features of different tabs and new evaluation metrics, TMWF achieved high-accuracy identification with fewer than six tabs. Deng et al. [109] introduced ARES, which is designed to perform WF attacks in multi-tab browsing scenarios. ARES formulates the multi-tab WF attack as a multi-label classification problem, where each classifier identifies whether a specific website is visited based on local traffic patterns. It uses a transformer-based model (Trans-WF) to extract and analyze local traffic patterns from multiple short traffic segments, even when the overall traffic pattern is obfuscated by WF defenses. In addition, ARES does not require prior knowledge of the number of open tabs, making it more practical for real-world deployment. Xie et al. [143] proposed a multi-scene WF method based on multi-head attention and data enhancement to address challenges in multi-tab and incomplete webpage traffic scenarios. For multi-tab browsing, the method (MAMWF) employs sequence embedding and block division to preserve traffic order, followed by multi-head attention to extract global features from mixed regions, improving classification accuracy. For incomplete traffic, the data enhancement-based webpage fingerprinting (DEWF) method uses generative adversarial imputation nets (GAIN) to enhance missing data, restoring fingerprint information for better recognition. Cui et al. [144] introduced Enhanced Deep Fingerprinting (EDF), a pair of DL-based WF attacks to overcome zero-delay lightweight defenses like FRONT and GLUE. The authors propose two models: EDF-S which integrates a global average pooling layer and a series structure into the existing Deep Fingerprinting (DF) framework, and EDF-P, which employs a parallel model for enhanced feature extraction.
@@
In 2024, Liu et al. [126] introduced FingerMamba, an approach to multi-tab WF using the Mamba-based structured state-space model (SSM). FingerMamba addresses the more realistic multi-tab scenario, where users open multiple tabs, increasing classification complexity. Yin et al. [13] investigated multi-tab website fingerprinting attacks (MTWFAs) using deep learning to analyze encrypted Tor traffic, focusing on realistic scenarios where users browse multiple websites simultaneously. The paper introduces MTWFA-SEG, an attack method that classifies each packet in multi-tab Tor traffic, enabling detailed inference of browsing activities and timing. The authors propose an improved fully convolutional network (FCN) model with skip transformations, residual connections, and a smoothing regularization technique to handle varying-length traffic traces effectively. Zhao et al. [33] presented Oscar, a WF attack framework designed to identify fine-grained webpages (e.g., subpages of the same website) rather than just websites, a challenge that traditional WF attacks struggle with. Existing WF attacks primarily classify entire websites, but their effectiveness drops significantly when trying to distinguish between different webpages due to high traffic similarity and large-scale dataset challenges. Unlike previous WF methods that focus on websites, Oscar applies metric learning to differentiate between similar-looking webpages by transforming the feature space. Chen et al. [158] presented RobustWF, a WF attack designed to handle multi-tab web browsing over encrypted tunnels. Traditional WF techniques struggle in realistic network environments due to dynamic tab concurrency, packet loss, duplication, and disorder. The proposed RobustWF method improves website fingerprinting accuracy and robustness using causality correlation and context learning. Instead of treating traffic as a mixed sequence, RobustWF groups packets into causality chains based on user requests and server responses. These chains help separate traffic from different websites, mitigating the tab concurrency challenge.
@@
In 2025, Meng et al. [132] presented Few-shot Multi-tab Website Fingerprinting (FMWF), an approach to multi-tab WF attacks that significantly reduces the need for large training datasets and frequent retraining. Unlike traditional WF models that require vast amounts of real-world multi-tab data, FMWF leverages transfer learning to adapt pre-trained models to multi-tab environments using only a small set of real traffic samples. Yuan et al. [119] introduced MW3F, a multi-tab WF attack that leverages Transformer-based feature fusion to enhance attack accuracy and robustness in encrypted networks like Tor. Traditional single-tab WF attacks are ineffective against multi-tab browsing, where traffic from different websites overlaps, making website identification difficult. MW3F overcomes this by integrating multi-head self-attention and learnable label embeddings to adaptively extract discriminative traffic features, even in the presence of strong WF defenses. Deng et al. [29] introduced CountMamba, a generalized WF attack framework that effectively identifies Tor users’ visited websites even in complex, real-world conditions. The CountMamba performs robustly across three key scenarios: robust attacks (resisting various defenses), early-stage attacks (working with partial traffic), and multi-tab attacks (handling concurrent browsing). Unlike traditional WF attacks that depend on full traffic traces and are limited to single-tab browsing, CountMamba combines Windowed Traffic Counting Matrix (WTCM) and State-Space-Oriented (SSO) Classifier. WTCM is a coarse-grained, time-window-based representation of traffic that counts packets based on direction, timing, and cell relationships. This method increases resilience against defenses that alter packet-level patterns. SSO is a fine-grained, causal, and iterative classifier that updates predictions as new traffic data arrives. This enables accurate early-stage identification without needing full traffic traces and provides insights into user activity across multiple tabs.

### v2

@@ V-A2 DL-based WF Attack models
Singh et al. [singh2023first] presented AutoKeras, a neural architecture search (NAS) tool, to explore whether automated DNN architecture design can improve WF attacks compared to manually crafted models. While NAS-generated models require larger exploration budgets and longer inference times, they highlight the potential for automated deep-learning-driven WF attacks. Zhuo et al. [zhou2023wf] introduced WF-Transformer, the first WF attack model based on Transformer networks, which is designed to capture the long-term dependencies in traffic traces, a feature that traditional CNN-based models fail to extract. The model uses packet direction as input and employs a sophisticated design of the Transformer network, including relative position embedding and multi-head attention mechanisms, to improve feature extraction. Jansen et al. [jansen2023repositioning] investigated the real-world feasibility of WF attacks on Tor by addressing a critical gap in previous research: the mismatch between training data collected from Tor exit relays and real-world WF attacks conducted from entry relays. To bridge this gap, the authors introduce Retracer, which transforms exit-side traces into entry-side traces, improving classifier training and yielding more accurate real-world WF threat estimates.
@@
To address the limitation associated with WF attacks and defense strategies in the presence of noise, Huang et al. [huang2023efficient] proposed a filter-assisted attack (FAA) to counter random packet defenses (RPD) and introduced a list-assisted defense (LAD) to mitigate FAA, balancing defense efficacy and overhead. The next year, Qian et al. [qian2024enhancing] investigated the impact of noise on WF attacks and introduced two attack methods: filter-assisted attack and augmentation-assisted attack. The first attack method leverages the packet size distribution to effectively filter out noise, while the second one trains a classification model by incorporating artificial noise. Experimental results show the efficacy of defending against different attacks and noises. Bahramali et al. [bahramali2023realistic] identified that WF classifiers often fail when deployed in unobserved network conditions due to variations in bandwidth, Tor circuits, and traffic patterns. To overcome this limitation, they proposed a network trace augmentation technique, NetAugment, to enhance the robustness of WF models. The NetAugment was instantiated through semi-supervised and self-supervised learning techniques.
@@
Through analyzing the present WF attacks, Mathews et al. [mathews2024laserbeak] found that a multi-channel input format that provides richer contextual information and enables the model to learn robust representations even in the presence of heavy traffic obfuscation. Then, the authors introduced Laserbeak, a state-of-the-art WF attack designed to break Tor defenses by leveraging multi-channel feature representation and transformer-based deep learning models. LASERBEAK demonstrates absolute performance improvements of up to 36.2% compared with prior attacks against defended traffic.
@@
The deep learning based WF attacks often ignore low-level timing information and use a simple direction-based representation method that only retains the spatial information of website traffic packets. However, the timing information plays a significant role in the WF attack under the research of Hong et al. [hong2024website]. They presented a WF attack technique called Time-Sampling, which leverages both spatial and temporal timing information from website traffic to improve the accuracy of WF attacks. The Time-Sampling model uses a time-based sampling technique that fills packet gaps with zeros, preserving both timing and spatial information in the traffic. Through evaluation, the accuracy is improved by 20% at most, higher than direction-based attacks.
@@
Traditional WF techniques either rely on single-flow features, which are sensitive to network changes, or multi-flow aggregation methods that fail to capture temporal and spatial relationships between network flows. To solve this issue, Tan et al. [tan2024inter] presented STC-WF, a WF attack that leverages spatio-temporal correlations among network flows to improve classification accuracy, particularly in encrypted traffic scenarios. The STC-WF uses a Graph Neural Network (GNN)-based approach that models these correlations as a Spatio-Temporal Correlation Graph (STCG). To overcome the limitation of traditional WF techniques, which struggle with content updates and labeling overhead, limiting their effectiveness in real-world scenarios. Gao et al. [gao2024multi] introduced MRCGCN, a self-supervised learning-based WF attack that leverages multi-level resource-coherent GNNs to improve the accuracy and robustness of WF attacks. MRCGCN overcomes these challenges by representing website traffic as structured graphs at resource, flow, and host levels and applying graph learning techniques for classification.
@@
Concept drift in website fingerprinting can occur when websites modify their structure or content. Even small changes, such as adding images, updating ads, or reorganizing page elements, alter the resulting traffic patterns. These shifts cause the distribution of features like packet sizes, timing, and traffic length to deviate from the data on which models were trained. To address the problem of concept drift, Ding et al. [ding2024multi] presented LRCT, a DL-based WF attack model designed to improve classification accuracy and efficiency while addressing concept drift—a phenomenon where model performance degrades over time due to network changes. The LRCT network effectively leverages the temporal learning advantages of Local RNN and the spatial learning strengths of CNN by designing the local feature extraction block (denoted as LRC Block). The network then uses a Transformer Encoder to capture more robust global features from the low-dimensional data. Experimental results comparing LRCT with state-of-the-art WF methods demonstrate that the proposed LRCT achieves 99.34% accuracy in a closed-world scenario, outperforming other benchmark models.
@@
Ginige et al. [ginige2024trafficgpt] proposed TrafficGPT, the first solution that leverages GPT-2 for feature extraction in encrypted traffic classification. GPT-2’s ability to understand sequential data is utilized to extract comprehensive features, improving the performance of open-set classification. This paper focuses on the open-set scenario, where the model must distinguish between known and unknown classes. This is particularly important in real-world applications, such as detecting harmful content while allowing benign traffic. Yuan et al. [yuan2025class] introduced Class Incremental Fingerprinting (CIF), a WF attack designed to continuously learn and expand its classification capabilities in real-world dynamic environments. Traditional DL-based WF attacks achieve high accuracy in static experimental settings but struggle in open-world scenarios, where new websites constantly emerge. CIF addresses this limitation by incorporating Class Incremental Learning (CIL) techniques, allowing the attack model to adapt to new monitored websites without forgetting previously learned knowledge. Yang et al. [yang2025dtpn] introduced DTPN (Diffusion-based Traffic Purification Network), which counteract adversarial perturbations in Tor WF defenses by leveraging diffusion models. Unlike traditional adversarial training, which struggles with generalization and frequent updates, DTPN purifies protected network traffic by removing adversarial noise, enabling existing WF classifiers to maintain high accuracy without retraining. The proposed UNetFormer architecture, combining TransformerBlocks and one-dimensional convolutions, estimates noise effectively, while a custom loss function enhances model generalization. Mei et al. [mei2025high] presented HPETC, a High Precision and Efficient Traffic Classification system designed to classify Tor traffic in real-world, large-scale network environments. Unlike previous Traffic Classification (TC) and WF techniques, HPETC effectively mitigates the base rate fallacy, a common issue where false positives overwhelm the classification process due to the extremely low proportion of anonymous traffic compared to normal network traffic.
@@ V-A3 Multi-tab WF attacks
BAPM [guan2021bapm],
Walkie-Talkie [wang2017walkie]
@@
ARES [deng2023robust] 91.30 91.30 90.20 81.80 75.20 58.70 52.80 43.80
FingerMamba [liu2024fingermamba]
@@
MTWFAs [yin2024traces] FCN
DS-19 [gong2020zero],
ARES [deng2023robust]
@@
FMWF [meng2025beyond] Few-shotFine-tuning
ARES [deng2023robust]
@@
ARES [deng2023robust],
Deng et al. [deng2024robust],
@@
ARES [deng2023robust] and
TMWF [jin2023transformer]
@@
Traditional WF attacks are performed on a well-constructed dataset, in which each packet trace is generated by visiting only one page tab at a time. Thus, attackers only need to analyze traces entirely belonging to a specific website without any noise data, which is known as the single-tab assumption. This assumption makes WF attacks less difficult, but it is unrealistic: users may engage in other activities, such as instant messaging alongside web browsing, and similar behavior is entirely random  [wang2016realistically]. In addition, even if users only browse the web, they may open a new page tab before the last one finishes loading, a behavior known as multi-tab browsing [cui2019revisiting, gu2015novel, juarez2014critical], such as jumping to another page through a hyperlink. A comparison of representative multi-tab WF attacks with 10% overlapping proportions as shown in TABLE IX.
@@
After Juarez et al. [juarez2014critical] pointed out that multi-tab browsing behavior significantly degrades the performance of WF models under the single-tab assumption, Wang and Goldberg [wang2016realistically] attempted to transform multi-tab traces into single-tab traces by splitting packet sequences and processing them using existing single-tab WF classifiers. In  [wang2016realistically], multi-tab traffic traces are categorized into three fundamental types based on their time interval relationships: Positive-time, Zero-time, and Negative-time. Figure 6 illustrates these scenarios for 2-tab traffic traces: (1) Positive-time separated traces refer to instances where the user lingers on a web page for a period before navigating to the next one, resulting in a brief lull in network activity and a noticeable gap in traffic. This time gap is significant enough to be detected as a pause but still shorter than the threshold used in time-based traffic splitting. (2) Zero-time separated traces refer to scenarios where a user clicks a link on a web page that is still loading, interrupting the current page and immediately initiating requests for the next one. This results in a clear transition between two web pages without any noticeable time gap between them. (3) Negative-time separated traces refer to situations where the user is loading two web pages simultaneously in multiple tabs. In such cases, the appropriate split point is considered to be the moment the second page begins loading. This is the most challenging class to separate, as there is neither a noticeable time gap nor a distinct cell pattern signaling the transition. Nonetheless, machine learning techniques can still be used to identify the split by extracting informative features from the cell sequences.
@@
Wang and Goldberg [wang2016realistically] focused on identifying split points within multi-tab traffic to divide it into multiple different single-tab segments for classification. However, this segmentation process results in information loss in the overlapping portions of the traffic. Although the algorithm improves the accuracy of split-point detection, it can only predict the first webpage visited in a multi-tab browsing session. Moreover, as the number of tabs increases, accurately locating the split points becomes increasingly difficult [yuan2025mw3f]. Xu et al. [xu2018multi] introduced the BalanceCascade-XGBoost algorithm to identify the split point between two tabs and classify traffic according to the initial chunk of data. Later, Yin et al. [yin2021automated] improved the BalanceCascade-XGBoost algorithm by applying Split Point Identification and Chunk-Based Classification mechanism. Cui et al. [cui2019revisiting] utilized the Hidden Markov Model to segment continuous browsing traffic and conducted webpage recognition on each segment through K-NN, achieving an impressive 80% accuracy in identifying the split points. To determine the final category of the webpage, a majority vote was employed. However, it treats overlapping areas equally with other areas and fails to consider the relevance and importance of the overlapping area in determining the webpage category [xie2023multi].
@@
In 2020, Cui et al. [cui2020more] utilized a CNN model to differentiate between single-tab and multi-tab traces, and extracted a fixed number of packets from the beginning and end of 2-tab traces. They then used CNN, Long Short-Term Memory (LSTM), and Stacked Denoising Auto-Encoding (SDAE) models to classify the extracted pure trace segments, further improving the accuracy of WF. However, their approach is still constrained by a limited number of page configurations and relies on multi-tab traces containing sufficiently long and non-overlapping pure segments. Then, Guan et al. [guan2021bapm] proposed a block attention profiling model (BAPM), which focuses on overlapping webpage traffic. However, with the number of webpages increasing, BAPM faces the challenge of dynamically adapting the number of attention heads, which introduces instability to the network structure and leads to more parameters. In addition, BAPM only considers a single scenario of overlapping visits and cannot adapt to more complex browsing scenarios. Previous CNN-based WF attacks rely on a carefully curated dataset where each packet trace is created by loading only one webpage at a time. Under this single-tab assumption, attackers analyze clean, noise-free traces, each belonging exclusively to a single website. However, the assumptions of previous work are obviously unreal. Users may access multiple webpages in multiple tabs at the same time.
@@
Guan et al. [guan2021bapm] found that the existing multi-tab attacks fail to exploit the overlapping regions, which contain mixed data from adjacent page tabs. To address this limitation, they leveraged the CNN to generate a tab-aware representation on the whole packet trace, including overlapping areas, and split the representation into blocks to reduce the influence of mixed data under the traffic direction features. Using multi-head attention, the model groups blocks belonging to the same website, allowing it to identify multiple websites simultaneously from a global perspective. This approach eliminates the need for explicit feature engineering and adapts naturally to different degrees of overlap.
@@
Later, Jin et al. [jin2023transformer] introduced Transformer-based Multi-tab Website Fingerprinting (TMWF), an end-to-end deep learning model that leverages a Transformer architecture to improve multi-tab WF attacks. Unlike previous models, TMWF does not require prior knowledge of the number of tabs in a browsing session. It treats multi-tab WF recognition as a set prediction problem, similar to object detection models. By introducing queries for tabs to extract WF features of different tabs and new evaluation metrics, TMWF achieved high-accuracy identification with fewer than six tabs. Deng et al. [deng2023robust] introduced ARES, which is designed to perform WF attacks in multi-tab browsing scenarios. ARES formulates the multi-tab WF attack as a multi-label classification problem, where each classifier identifies whether a specific website is visited based on local traffic patterns. It uses a transformer-based model (Trans-WF) to extract and analyze local traffic patterns from multiple short traffic segments, even when the overall traffic pattern is obfuscated by WF defenses. In addition, ARES does not require prior knowledge of the number of open tabs, making it more practical for real-world deployment. Xie et al. [xie2023multi] proposed a multi-scene WF method based on multi-head attention and data enhancement to address challenges in multi-tab and incomplete webpage traffic scenarios. For multi-tab browsing, the method (MAMWF) employs sequence embedding and block division to preserve traffic order, followed by multi-head attention to extract global features from mixed regions, improving classification accuracy. For incomplete traffic, the data enhancement-based webpage fingerprinting (DEWF) method uses generative adversarial imputation nets (GAIN) to enhance missing data, restoring fingerprint information for better recognition. Cui et al. [cui2023edf] introduced Enhanced Deep Fingerprinting (EDF), a pair of DL-based WF attacks to overcome zero-delay lightweight defenses like FRONT and GLUE. The authors propose two models: EDF-S which integrates a global average pooling layer and a series structure into the existing Deep Fingerprinting (DF) framework, and EDF-P, which employs a parallel model for enhanced feature extraction.
@@
In 2024, Liu et al. [liu2024fingermamba] introduced FingerMamba, an approach to multi-tab WF using the Mamba-based structured state-space model (SSM). FingerMamba addresses the more realistic multi-tab scenario, where users open multiple tabs, increasing classification complexity. Yin et al. [yin2024traces] investigated multi-tab website fingerprinting attacks (MTWFAs) using deep learning to analyze encrypted Tor traffic, focusing on realistic scenarios where users browse multiple websites simultaneously. The paper introduces MTWFA-SEG, an attack method that classifies each packet in multi-tab Tor traffic, enabling detailed inference of browsing activities and timing. The authors propose an improved fully convolutional network (FCN) model with skip transformations, residual connections, and a smoothing regularization technique to handle varying-length traffic traces effectively. Zhao et al. [zhao2024towards] presented Oscar, a WF attack framework designed to identify fine-grained webpages (e.g., subpages of the same website) rather than just websites, a challenge that traditional WF attacks struggle with. Existing WF attacks primarily classify entire websites, but their effectiveness drops significantly when trying to distinguish between different webpages due to high traffic similarity and large-scale dataset challenges. Unlike previous WF methods that focus on websites, Oscar applies metric learning to differentiate between similar-looking webpages by transforming the feature space. Chen et al. [chen2024causality] presented RobustWF, a WF attack designed to handle multi-tab web browsing over encrypted tunnels. Traditional WF techniques struggle in realistic network environments due to dynamic tab concurrency, packet loss, duplication, and disorder. The proposed RobustWF method improves website fingerprinting accuracy and robustness using causality correlation and context learning. Instead of treating traffic as a mixed sequence, RobustWF groups packets into causality chains based on user requests and server responses. These chains help separate traffic from different websites, mitigating the tab concurrency challenge.
@@
In 2025, Meng et al. [meng2025beyond] presented Few-shot Multi-tab Website Fingerprinting (FMWF), an approach to multi-tab WF attacks that significantly reduces the need for large training datasets and frequent retraining. Unlike traditional WF models that require vast amounts of real-world multi-tab data, FMWF leverages transfer learning to adapt pre-trained models to multi-tab environments using only a small set of real traffic samples. Yuan et al. [yuan2025mw3f] introduced MW3F, a multi-tab WF attack that leverages Transformer-based feature fusion to enhance attack accuracy and robustness in encrypted networks like Tor. Traditional single-tab WF attacks are ineffective against multi-tab browsing, where traffic from different websites overlaps, making website identification difficult. MW3F overcomes this by integrating multi-head self-attention and learnable label embeddings to adaptively extract discriminative traffic features, even in the presence of strong WF defenses. Deng et al. [deng2025countmamba] introduced CountMamba, a generalized WF attack framework that effectively identifies Tor users’ visited websites even in complex, real-world conditions. The CountMamba performs robustly across three key scenarios: robust attacks (resisting various defenses), early-stage attacks (working with partial traffic), and multi-tab attacks (handling concurrent browsing). Unlike traditional WF attacks that depend on full traffic traces and are limited to single-tab browsing, CountMamba combines Windowed Traffic Counting Matrix (WTCM) and State-Space-Oriented (SSO) Classifier. WTCM is a coarse-grained, time-window-based representation of traffic that counts packets based on direction, timing, and cell relationships. This method increases resilience against defenses that alter packet-level patterns. SSO is a fine-grained, causal, and iterative classifier that updates predictions as new traffic data arrives. This enables accurate early-stage identification without needing full traffic traces and provides insights into user activity across multiple tabs.